OpenStudy (anonymous):

Simplify m - {n - [p + (m + n - p)]}. 0 2m -2m + 2p

OpenStudy (anonymous):

First distribute the first negative sign and remove the {}. m - n + [p + (m + n - p)] Distribute the positive sign before the [ and remove the [] m - n + p + (m + n - p) Distribute the positive sign before ( and remove the () m - n + p + m + n - p Group like terms, i.e. (m with m , n with n and p with p). (m + m) + (n - n) + (p - p) = 2m + 0 + 0 = 2m hope that helps! :)!

OpenStudy (anonymous):

yes it does tthanks

OpenStudy (anonymous):

or m - {n - [p + (m + n - p)]} = m -{n - [p + m + n - p]} = m - {n - p - m - n + p} = m - n + p + m + n - p = 2m